In Desmos, using the Rational Root Theorem type the equation 2x3+ 5x2- 4 -3


which of the following match the possible roots that are NOT the actual roots of the function?


1
Expert's answer
2021-04-29T14:19:29-0400

"2x^3+ 5x^2- 4x -3 \\
p= ± 1, ± 3 \\ q= ± 1, ± 2 \\
Possible \ rational \ roots \ are: \ x=p/q=±\{1,3,\frac{1}{2},\frac{3}{2} \} \\ 
Answer: \ x=-1, \ 3, \ \frac{1}{2}, \ \frac{3}{2}, -\frac{3}{2}"